Reported the following data for 2012.Labour force.17948 000.Employment l, 13422000,working age population 32786 000
Calculate the unemployment rate
1
Expert's answer
2021-09-17T12:06:04-0400

Given here the data for the labor force of 2012:

Labor force: 17,948,000

Employment: 13,422,000

Working Age Population: 32,786,000


Hence, the following can be calculated. 

The unemployment rate can be defined as the part of the labour force that is unemployed but actively looking for a job over the last 30 days. It can be calculated with the following formula:

"Unemployment\space Rate =\frac{ Number \space of\space people\space Unemployed}{

Labour\space Force

}

×100"


Since the number of unemployed people is not given, it can be calculated as the part of the labor force that is not employed, with the help of the following formula:

"Labour \space Force = Employed+Unemployed"


Here, the labor force and the number of employed is given, which will help us calculate the number of unemployed.

"Labour \space Force = Employed+Unemployed\\17,984,000=13,422,000+Unemployed\\Unemployed=17,984,000-13,422,00\\=4,586,000"


Hence, the unemployment rate is:

"Unemployment \space Rate =\frac{ 4,596,000}{17,984,000}×100=25.5\%"

The unemployment rate is 25.5% in the economy.